OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1) 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = 4(4n+1)(8n+7)/6 2) 1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = n(6n^2-3n-1)/2

OpenStudy (helder_edwin):

i'll do the first one. and u can do the second one. OK?

OpenStudy (anonymous):

Okay I'll try :)

OpenStudy (helder_edwin):

let's try the second. it seems you have some typos in the first one. check that out.!!

OpenStudy (helder_edwin):

we have to prove for all n natural that \[ \large 1^2+4^2+7^2+\dots+(3n-2)^2=\frac{n(6n^2-3n-1)}{2} \]

OpenStudy (helder_edwin):

(i) n=1 (induction base). If n=1 the left side is \[ \large 1^2=1 \] and the right side is \[ \large \frac{1(6\cdot1^2-3\cdot1-1)}{2}=\frac{6-3-1}{2}=\frac{2}{2}=1 \] they are equal, hence we have our induction base. OK?

OpenStudy (anonymous):

Okay...

OpenStudy (helder_edwin):

Now, let n>1 and suppose that (induction hypothesis) \[ \large 1^2+4^2+7^2+\dots+(3n-2)^2=\frac{n(6n^2-3n-1)}{2} \] we have to prove that (induction thesis) \[ \large 1^2+\dots+(3n-2)^2+[3(n+1)-2]^2=\frac{(n+1)[6(n+1)^2-3(n+1)-1]}{2} \] OK?

OpenStudy (helder_edwin):

oh sorry the right side should be \[ \large =\frac{(n+1)[6(n+1)^2-3(n+1)-1]}{2} \]

OpenStudy (helder_edwin):

let's reduce a little bit: \[ \large =\frac{(n+1)[6(n^2+2n+1)-3n-3-1]}{2} \] \[ \large =\frac{(n+1)(6n^2+9n+2)}{2} \] so our induction thesis is \[ \large 1^2+\dots+(3n-2)^2+(3n+1)^2= \frac{(n+1)(6n^2+9n+2)}{2} \] r u following so far??

OpenStudy (helder_edwin):

@aleisha0301 ??

OpenStudy (anonymous):

Oh yes I am! Sorry

OpenStudy (helder_edwin):

Ok. let's probe the thesis: \[ \large 1^2+\dots+(3n-2)^2+(3n+1)^2= \] \[ \large =\frac{n(6n^2-3n-1)}{2}+(3n+1)^2 \] \[ \large =\frac{n(6n^2-3n-1)+2(9n^2+6n+1)}{2} \] \[ \large =\frac{6n^3-3n^2-n+18n^2+12n+2}{2} \] \[ \large =\frac{6n^3+15n^2+11n+2}{2} \] ok so far?

OpenStudy (anonymous):

Yes I understand so far :)

OpenStudy (helder_edwin):

we r gonna use synthetic division to factor the last numerator into what we r supposed to get: \[ \large \begin{array}{r|rrrr} & 6 & 15 & 11 & 2\\ -1 & & -6 & -9 & -2 \\ \hline & 6 & 9 & 2 & \fbox{0} \end{array} \] this gives us \[ \large =\frac{6n^3+15n^2+11n+2}{2}= \frac{(n+1)(6n^2+9n+2)}{2} \] which is precisely what we were supposed to get. QED.

OpenStudy (helder_edwin):

@aleisha0301 ?? did u follow all of this??

OpenStudy (anonymous):

I did. So the last part is the answer?

OpenStudy (helder_edwin):

everything is the answer. it is a demonstration by induction.

OpenStudy (anonymous):

Hm okay. Thank you for taking time to help me! :)

OpenStudy (helder_edwin):

glad to help.

OpenStudy (helder_edwin):

now for the first one, there is something wrong. u have \[ \large 4\cdot6+5\cdot7+6\cdot8+\dots+4n(4n+2)= \frac{4(4n+1)(8n+7)}{6} \]

OpenStudy (helder_edwin):

do u see the last summand on the left side.??

OpenStudy (helder_edwin):

namely 4n(4n+2)

OpenStudy (helder_edwin):

@aleisha0301 ??

OpenStudy (anonymous):

In the problem it's 4(4n+1)(8n+7)/6

OpenStudy (helder_edwin):

no, on the left side. ignore the right side.

OpenStudy (anonymous):

Ohh! it's 4(4n+2). My mistake!

OpenStudy (helder_edwin):

why don't repost the problem (properly written) and will work on it. OK??

OpenStudy (anonymous):

Okay :)

OpenStudy (helder_edwin):

just "call" me.

OpenStudy (anonymous):

I want to say thank you this help me understand so much

OpenStudy (ally_b):

so the statement is true for all positive integers? @helder_edwin

OpenStudy (ally_b):

@aleisha0301

OpenStudy (helder_edwin):

I don't know, @Ally_B . She never posted the problem with the corrections.